Write the equation of the bisector of the segment joining the points with coordinates of (-3,4) and (5,-2)
thank you

Answer by stanbon(75887)

Write the equation of the bisector of the segment joining the points with coordinates of (-3,4) and (5,-2)
-----
Plot the points so you can see the solution.
---
Find the midpoint (x,y) of the line segment joining the points.
x = (-3+5)/2 = 1
y = (4-2)/2 = 1
-----
mid-point is (1,1)
========================
Find the slope of the line segment:
m = (-2-4)/(5--2) = -6/7
=============================
Comment: I'm going to assume you want the perpendicular bisector.
Fact: The pb must have a slope of 7/6 and must pass thru (1,1)
=============================
Form: y = mx+b
1 = (7/6)(1) +b
b = -(1/6)
----
Equation:
y = (7/6)x-(1/6)
========================
cheers,
Stan H.